Use the t distribution to find the reliability factor for a confidence interval based on the following

confidence coefficients=0.99 and sample sizes: 24


1
Expert's answer
2021-11-11T14:44:13-0500

The confidence coefficientis 0.99

(i.e) 1-α=0.99 ---> α=0.01

"\frac{α}{2}=\frac{0.01}{2}=0.005"

Sample size, n=24

Degrees of freedom=n-1=23

The t – table value with 23d.f is

"t_ {\u03b1\/2}=2.807" (Fromt tables)